If Maggie has $5,000 to invest and wants to have $10,000 at the end of 9 years, what interest rate must she earn on her money, assuming annual compounding?

Select one:

a. 5%

b. 6%

c. 7%

d. 8%

e. None of the above

Please show all work and use present value tables for caculations
Answer #1
FV =PV(1+i)^n $10,000 =$5,000*(1+i)^9
(1+i)^9 =$10,000/$5,000
(1+i)^9 =2
I =8%
You can also use the PV table for calculation
PV of 8% for 9 years which is PVIF(8%,9 years) is 0.50025
So FV =$10,000
PV =$10,000*0.50025 =$5,000(apprx)
So Option D is answer